# Dynamical phases of short-term memory mechanisms in RNNs

### v1

@@ Acknowledgements
  • Unstable Training: Training is classified as unstable if, after achieving accuracy and reliability values greater than $0.8$, the performance falls below $0.6$ for more than $300$ epochs.
  • Failure to Converge to a Local Minimum: An RNN is considered to have failed to converge to a local minimum if the mean accuracy and reliability over the last $100$ epochs are less than $0.7$.
  • Limit Cycle: An RNN is classified as exhibiting a limit cycle if the accuracy and reliability values for the last $100$ epochs are $\geq 0.8$ and the oscillation measure (OM) is $\geq 0.1$.
  • Ghost Point: An RNN is classified as exhibiting a ghost point if the accuracy and reliability values for the last $100$ epochs are $\geq 0.8$ and the oscillation measure (OM) is $<0.1$.
Jointly, the first two classes made up the “not learned” classification.
S2.3 Estimating Cutoff Delay Values
To estimate the cutoff delay values for each learning rate, we fitted a modified ReLU function $f(x)$ with learnable parameters $a$ and $b$ to the ”not learned” distribution of the corresponding learning rate $X\in\mathbb{R}^{N}$, as shown in Figure 6 (C-D).
$f(x)=\begin{cases}bx&\text{if }x>a\\ 0&\text{if }x\leq a\end{cases}$ (S33)
We define the loss function $\mathcal{L}(a,b)$ as follows:
$\min_{a,b}\mathcal{L}(a,b)=\sum_{i=1}^{N}\left(f(x_{i};a,b)-y_{i}\right)^{2}$ (S34)
Upon completion of the optimization process, the parameter $a$ is identified as the cutoff delay value for the corresponding learning rate.
Appendix S3 Supplementary figures
Figure S1: An example set of training curves for the large scale RNN experiment with no post-reaction period. The plots show the training results reported in Fig. 6C, E for the learning rate value of $10^{-2}$.
Figure S2: The phase diagram for the RNNs trained to perform cue-associated responses. Similar to the delayed cue discrimination tasks, we trained RNNs to provide outputs without any delays. We varied the length of the response (after the cue window) and enforced the networks to output one with the onset of the cue. The ability to learn the task was only mildly dependent on the response window length (compare to Fig. 7 and note the orders of magnitude larger differences in learning rates), indicating that the short-term memory component of the delayed cue discrimination task is primarily responsible for the scaling law. Parameters: $T_{\rm in}=30ms$, $\tau=10ms$, $\Delta t=5ms$, learning rates and the lengths of the response windows are varied. In total, we trained 8,000 RNNs on this task ($20$ different response windows and $20$ learning rates, $20$ networks per condition).

### v2

@@ Acknowledgements
  • Unstable Training: Training is classified as unstable if, after initially achieving accuracy and reliability values above $0.8$, the performance subsequently falls below $0.6$ for more than 10% of the epochs following the initial success (Fig. S4B).
  • Limit Cycle: An RNN is classified as exhibiting a limit cycle if its mean accuracy and reliability during the final 10% of epochs are $\geq 0.8$, and the Mechanism Discrimination Index (MDI) is less than $0.5$ (Fig. S4C).
  • Ghost Point: An RNN is classified as exhibiting a ghost point if its mean accuracy and reliability during the final 10% of epochs are $\geq 0.8$, and the Mechanism Discrimination Index (MDI) is greater than or equal to $0.5$ (Fig. S4D).
The first two categories collectively constitute the “Not Learned” classification.
S3.4.2 Plotting the scaling laws
To investigate the scaling relationship between the learnable delay interval and the learning rate, we performed the following analysis.
First, for each learning rate, we extracted the model’s performance across different delay intervals from the large-scale experiments (Fig. 4B, E). We then fit a saturating function—either an exponential saturation or a sigmoid—to the performance curves over delay (Fig. S3D, E). This yielded a cutoff delay value $a^{*}$, representing the point at which performance saturates or exhibits a sharp transition. Specifically, the exponential saturation fit was defined as:
$f(x;x_{0},\alpha)=\begin{cases}0&\text{if }x<x_{0}\\ 1-\exp(-\alpha(x-x_{0}))&\text{otherwise}\end{cases}$ (S39)
The fits were applied to performance metrics (e.g., reaction accuracy or reliability) over delay intervals $x$, using scipy.optimize.curve_fit.
Next, we analyzed how the cutoff delay values $a^{*}$ scale with learning rate $\alpha$. We performed a log-log linear regression of the form:
$\log(\alpha)=\beta\log(a^{*})+\log(A)$ (S40)
The critical exponent $\beta$ was estimated using scipy.stats.linregress and validated by an additional log-log fit via curve_fit. To estimate confidence intervals, we performed Monte Carlo sampling using the parameter covariance matrix. The resulting fit and its 95% confidence interval were visualized on log-log axes, as shown in Fig. 4C, F and Fig. S3F.
This analysis enabled us to extract empirical scaling laws for slow-point manifolds and limit-cycles, which were then compared to theoretical predictions.
S3.4.3 Plotting Firing Rates
To visualize firing rates, we collect the hidden states of the leaky firing-rate RNN during inference. To reveal the underlying dynamical structure, we extend the trial by appending a post-reaction interval ($T_{\rm post}$) that is nine times longer than the original duration, defined as $T=T_{\rm in}+T_{\rm delay}+T_{\rm resp}$.
After collecting the full firing rate trajectory, we sort neurons based on the timing of the peak activation in their absolute firing rates. Specifically, each neuron’s activity is scanned to identify the time step at which its firing rate reaches its maximum value. Neurons that do not exceed a predefined activation threshold are pushed to the end of the ordering. This procedure ensures that neurons are visualized in the order in which they become active, highlighting sequential structure in the network’s dynamics.
Appendix S4 Reproduction details
Figure 2:
To investigate the emergent mechanisms—specifically, slow-point manifolds and limit cycles—we conducted experiments using low-rank recurrent neural networks (RNNs) (Background). Given that a minimum of two dimensions is required for the formation of limit cycles, we set the low-rank dimensionality to $K=2$. The delayed activation task was chosen as the training paradigm. The learning rate was initialized at $\alpha=10^{-2}$. Additionally, we set the time decay constant to $\tau=10$ ms, the time step to $\Delta t=5$ ms, the delay interval to $T_{\rm delay}=150$ ms, and the response interval to $T_{\rm resp}=50$ ms. The architecture was trained for $150,000$ epochs with a total of $N=100$ neurons. All weights and biases, as described in Eq. (2), were optimized throughout training. For this experiment, noise was set to $\epsilon=0$, and firing rates were initialized from a multivariate Gaussian distribution with mean $\mu_{r}=0$ and standard deviation $\sigma_{r}=0.1$. Training was performed using stochastic gradient descent (SGD) with PyTorch’s default hyperparameters.
Figure 3:
This figure highlights the impact of even slight changes in task configuration on network dynamics. To demonstrate this, we trained identical RNNs with $N=100$ neurons and a rank of $K=2$, by increasing the post-reaction period interval from $T_{\rm post}=0$ ms to $T_{\rm post}=30$ ms. We applied a similar configuration to Fig. 2A, except for setting $T_{\rm delay}=30$ ms, $T_{\rm resp}=10$ ms, and $\alpha=10^{-3}$ for the experiments in Fig. S1A-B, training the networks for $10^{6}$ epochs. For the experiments in Fig. S1C, we set $\alpha=5\times 10^{-3}$ and trained the networks for $500,000$ epochs while varying the delay interval as $T_{\rm delay}\in\{30,40,50,60,70\}$ ms. Moreover, SGD is used with the default hyperparameters of PyTorch.
Figure 4:
In large-scale RNN experiments, we developed a primary training pipeline to systematically investigate the effects of learning rate, delay interval length, and post-reaction interval. Apart from these variations, all other training configurations in Fig. 4 were set as follows: number of neurons $N=100$, noise initialization drawn from a Gaussian distribution with mean $\mu_{\epsilon}=0$ and standard deviation $\sigma_{\epsilon}=10^{-3}$, and firing rate initialization drawn from a Gaussian distribution with mean $\mu_{r}=0$ and standard deviation $\sigma_{r}=10^{-1}$. Additionally, task-related parameters were fixed, with the cue interval set to $T_{\rm in}=30$ ms, the reaction interval to $T_{\rm resp}=50$ ms, the time decay constant to $\tau=10$ ms, and the time step to $\Delta t=5$ ms. Since lower learning rates require longer training durations, we implemented an adaptive schedule for the number of training epochs based on the learning rate. The number of epochs was determined using the following equation:
[…]
The delay interval was gradually increased in increments of $20$ ms, covering the range $T_{\rm delay}\in[0,420]$ ms. In our initial experiments, learning rates were sampled from a logarithmic scale with equal spacing in the range $\alpha\in[10^{-3},10^{-1}]$. To further investigate the effects of intermediate learning rates, we additionally sampled more values logarithmically within the range $\alpha\in[10^{-1.6},10^{-0.5}]$ (see Table S1 for details). Besides, Fig. 4C,F presents the estimated cutoff delay values, which are described in detail in Section 2.5. Moreover, in Fig. H, I, we constructed a phase diagram of emergent mechanistic phases derived from the results shown in Fig. 4C-D.
Figure S2:
We analyze the emergence of the limit cycle model in the delayed activation task through analytical methods (Supplementary 1.2). Figure S2A illustrates the vector field of the dynamical system, which is derived from Eq. (S15) with $r=0.1$. Additionally, Figure 4G presents two loss curves: the analytical curve is obtained from Eq. (S17), while the numerical curve is computed using Eq. (S20), with parameters set to $T_{\rm delay}=60$ ms, $T_{\rm resp}=10$ ms, and $dt=10^{-2}$. Finally, Figure S2B is constructed by plotting analytically derived scaling laws, with the response interval set to $T_{\rm resp}=10$ ms.
Figure S3:
This figure builds on the results presented in Fig. 4 to illustrate the RA and MDI metrics. Additionally, it highlights the scaling laws associated with slow-point manifolds observed in the large-scale RNN experiments.
Figure S4:
We randomly sampled the example classes of dynamical systems from the large-scale RNN experiments.
Figure S5:
In this experiment, we use the same training configuration as in Fig. 4, except for the delay intervals and the value of $\tau$. We set $\tau=20$ ms and sampled $T_{\rm delay}$ from the range $[0,840]$ ms in increments of 40 ms. (See Table S1 for more details.)
Figure S6:
In this figure, the only differences from the large-scale experiments are that the number of training epochs is set to $3000$, $T_{\rm delay}=0$ ms, and the response interval $T_{\rm resp}$ is varied instead. This setup allows us to examine the importance of the memory component. (See Table S1 for more details.)
The code to reproduce these figures is publicly available at https://github.com/fatihdinc/dynamical-phases-stm, the data jointly with the code at https://doi.org/10.5281/zenodo.15529757.
Figures and captions
Figure S1: RNNs can switch between distinct short-term memory mechanisms during training. A To illustrate that different mechanisms can be employed by dynamical systems to delay their responses, we trained a rank-2 RNN to perform the delayed activation task. In this task, networks initialized at a given state (Methods) should delay their output by $T_{\rm delay}$. Parameters: $N=100$ neurons, $\tau=10$ms, $\Delta t=5$ms, $T_{\rm delay}=150$ms, $T_{\rm resp}=50$ms, $\alpha=10^{-2}$ for stochastic gradient descent using otherwise the default parameters in PyTorch (Paszke et al., 2017). The dotted lines correspond to epochs, for which the latent dynamical systems are illustrated in Panel (B). B During the training, two distinct mechanisms emerge in the RNN, which we probe by plotting the two-dimensional latent dynamical systems. Left. Early in the training, the network lowers the loss values by designing a line of slow-points (referred to as SP manifold), towards which other stats are attracted. The latent states after the trial concludes end up in a final attractive fixed-point. Middle. As training progresses, the SP manifold slowly dissolves into a circular pattern, which we deem as a transient mechanism before the network settles into the final solution. Right. After further training, a limit cycle emerges in the latent dynamical system. In this structure, even after the trial concludes, latent states continue to evolve in a periodic manner. The shaded areas around the decision boundary correspond to network outputs within $0.75$ (outer green boundary) and $0.25$ (outer red boundary).
Figure S2: A theoretical scaling analysis of short-term memory mechanisms via toy models. A The illustration of the flow map corresponding to the toy limit cycle model for $r=0.1$. The red circle denotes the attractive limit cycle, the purple line corresponds to the decision boundary of the output, and the green circle denotes the starting point of the state. B An example loss curve for the limit cycle toy model, with analytical curve and empirical data points. Note the kink at the global minimum. C The different dependencies of the effective learning rate for two types of mechanisms create a phase space as a function of delay and learning rate values. In general, there are three main phases of short-term memory mechanisms for large $T_{\rm delay}\gg T_{\rm resp}$. For this figure, we picked $T_{\rm resp}=10$ms.
Figure S3: Details of the large-scale RNN experiment. A Different memory mechanisms exhibit distinct loss curves during training. While limit cycles and slow-point manifolds successfully converge to local minima, the “Final Model Failure” and “Unstable Learning” classes do not achieve stable convergence due to high learning rates. B As further illustrated, only the limit cycle and slow-point manifold mechanisms achieve consistently high reaction accuracies. C The Mechanism Discrimination Index (MDI) (see Methods) allows reliable differentiation between limit cycles and slow-point manifolds, even though their dynamics appear similar during the trial period. D Lowering the learning rate enhances the model’s ability to learn tasks with longer delay intervals. The solid curve represents the corresponding exponential saturation fit (see Methods). E Additionally, decreasing the learning rate increases the likelihood of models adopting slow-point manifolds. F As predicted by analytical models, results from the large-scale RNN experiments exhibit a similar power-law relationship ($T_{\rm delay}^{-4}\geq T_{\rm delay}^{\beta}\geq T_{\rm delay}^{-5}$) for learning slow-point manifolds for large learning rates. However, for larger delays ($T_{\rm delay}>100ms$), these solutions preferentially suppressed, presumably in favor of the limit cycle based solutions.
Figure S4: The four main classes identified in the large-scale RNN experiment. This figure illustrates representative examples from each of the four primary classes: A Final model failure, B Unstable learning, C Limit cycle, and D Slow-point manifold. We focused on trained models from experiments with a delay interval of $180$ ms, which was specifically selected because it exhibits all four classes, as illustrated in Fig. 4B. The first column presents the reaction metrics, i.e., Reaction Accuracy and Reaction Reliability (see Methods). The second column shows the network’s outputs alongside the corresponding ground truth signals. The third column displays the firing rates of individual neurons. The final column illustrates the first two principal components derived from these firing rates. Training parameters used for each representative example are as follows: A $\alpha=0.1$, seed = $93$; B $\alpha\approx 0.01931$, seed = $41$; C $\alpha=0.1$, seed = $82$; and D $\alpha=0.01$, seed = $85$. Additional parameters shared across all models are $T_{\rm in}=30$ ms, $T_{\rm delay}=180$ ms, $T_{resp}=50$ ms, $\Delta t=5$ ms, $\tau=10$ ms, $\sigma_{\rm noise}=0.001$, $\sigma_{\rm r}=0.1$, and batch size per channel = $64$ (see Methods).
Figure S5: RNN experiments with increased $\tau$. In this figure, we examine the effect of increasing the time constant to $\tau=20$ ms while keeping all other hyperparameters identical to those used in the large-scale delayed cue-discrimination task experiments (Fig. 4). These experiments reveal that increasing $\tau$ does not alter the observed scaling laws. A As in Fig. S3, we observe similar loss curve patterns across all classes under the larger $\tau$ setting. B Consistent with previous results, the distribution of dynamical classes across different random seeds remains largely unchanged with the increased $\tau$. C The phase diagram for $\tau=20$ ms shows scaling laws consistent with those observed in the smaller $\tau$ regime.
Figure S6: Scaling laws disappear in fixed-response task experiments with RNNs. In this experiment, we remove the delay period from the delayed cue-discrimination task to isolate the effect of memory and assess its role in scaling laws. As shown in the figure, the removal of the delay component eliminates the previously observed scaling behavior, highlighting memory as a critical factor underlying these laws. A As in previous tasks, we observe qualitatively similar loss curves between not learned and correct networks. B Phase diagram showing the classification of not learned and correct networks in the fixed-response task. The x-axis represents the length of the response interval, during which the model is required to generate a consistent output over time. Notably, removing the delay period (and thus the memory requirement) leads to a substantial loss of scaling structure. C We compute the scaling law using the same methodology as in Fig. 4C,F and Fig. S3F.